ops2 keeps freezing and is just generally slow, I could not even open gedit for the summary files (had to load them on ops4) (Ellen)

Hobart 12m:

Disk VSN: HAY-0047

Data volume at beginning: O GB

Katherine 12m:

Disk VSN: HOB+0041

Data volume at beginning: 0 GB

  • Jump in clkoff time since experiment setup 10 hours earlier, but appears to be stable now (Ross).
  • 18:15:19 UT: Large difference between formatter and maser delays (Ross).
  • 18:30 UT: Field system not responding, schedule could not be restarted after fmset (Ross).
  • 19:00 UT: Schedule restarted and time difference stable (Ross).
  • 19:07 UT: First reliable scan 316-1907, 54.84 GB (Ross).
  • 03:21 UT: clock difference has drifted over the last two hours. Alarm of log monitor @ Mt Pleasant not working. DBBC reconfigured. Still drifting. Power cycle the DBBC finally helped. Scans missed between 3:30 and 4:00. Missed scans 317-0332 to 317-0349. (Lucia)

Yarragadee 12m:

  • 15:25 UT: Autocorrelations show distinct signals. Jamie says we see the phase cal tones here. (Lucia)

Disk VSN: HOB+0022

Data volume at beginning: 0 GB